*** Your title here ***

*** Your name here *** **City**, 2014

Vilnius University Institute of Biotechnology

1 / 4

*** Slide title ***

...

[Gražulis et al., 2009 ]

2 / 4

Thank you!

http://en.wikipedia.org/wiki/Topaz  http://www.crystallography.net/2207377.html

References

Gražulis, S., Chateigner, D., Downs, R. T., Yokochi, A. F. T., Quirós, M., Lutterotti, L., Manakova, E., Butkus, J., Moeck, P., and Le Bail, A. (2009). Crystallography Open Database – an open-access collection of crystal structures. Journal of Applied Crystallography, 42:726–729.